### v1

@@ Abstract
Figure 1: Overview of contributions: (a) WorkArena is a benchmark of 29 web tasks and 23,150 unique instances that cover common ways of interacting with the ServiceNow Platform, a widely-used enterprise software platform. (b) BrowserGym is a Python environment for designing and evaluating web agents, which includes a rich set of actions and multimodal observations (shown here the HTML contents of the page, its accessibility tree, and the raw pixels after browser rendering).
@@ 1 Introduction
  • WorkArena: A realistic benchmark of enterprise-related tasks for web agents comprising 23,150 unique task instances (§ 3, Fig. 1a);
  • BrowserGym: A new environment for the development and evaluation of web agents, compatible with previous benchmarks like WebArena (Zhou et al., 2023), MiniWoB (Liu et al., 2018; Shi et al., 2017a) and WebShop (Yao et al., 2022), that offers a richer set of multimodal observations (e.g., screenshot, accessibility tree, screen coordinates), a broader set of actions (e.g., Python code and high-level primitives), and supports chat-based interactions (§ 4, Fig. 1b);
@@ 2 Related Works
Benchmarks for web agents:  Early benchmarks for web agents were based on synthetic web environments where agents were tasked with performing low-level keyboard and mouse actions (Shi et al., 2017b). Notable examples are MiniWoB (Shi et al., 2017a; Liu et al., 2018), which offer a collection of 125 toy web tasks ranging from clicking a specific button to using a basic text editor, and WebShop (Yao et al., 2022), a simulated e-commerce website with shopping tasks that require searching and browsing a catalog of items. More recently, Zhou et al. (2023) introduced WebArena, a collection of 190 tasks based on realistic websites that emulate real-world domains such as e-commerce, social forums, collaborative software development, and content management. WebArena is a notoriously challenging benchmark, with a success rate of 14% for a state-of-the-art web agent based on GPT-4, and 78% for human agents. Deng et al. (2023) proposed Min2Web, a large-scale dataset of 2,000 web interactions from 137 website curated by human annotators. Last, He et al. (2024) propose 300 information-retrieval tasks, from 15 real-world consumer websites (e.g., Amazon, Coursera, Booking), which are used to evaluate WebVoyager, a vision-based web agent.
@@ 3 WorkArena – An Enterprise Benchmark
WorkArena consists of a suite of 29 tasks and 23,150 unique instances that cover core interactions with the ServiceNow platform, such as navigating lists, filling forms, searching knowledge bases, utilizing service catalogs, and navigating via menus (see Fig. 0(a)). Collectively, these tasks are representative of a wide array of common operations that employees, like IT, administrative, and white-collar staff, perform on a daily basis.
@@ 3.1 WorkArena Tasks
Lists:  We consider 12 list-based tasks, which can be grouped into two categories: filtering and sorting. The former consists of using the UI to construct a complex filter with 1 to 5 conditions. The latter consists of using the UI to sort the list based on up to 3 columns. In both cases, the interaction with the UI is non-trivial and requires opening a hidden menu, adding the right number of conditions, and filling them out accordingly. There are 6 tasks of each type, each corresponding to a different data table (e.g., users, incidents). In both cases, client-side validation is used to verify that the resulting list satisfies the expected conditions. Together, these tasks yield 12,000 instances.
@@ 4.2 An Ideal Experimental Framework
Extensibility:  BrowserGym is easily extensible to additional benchmarks. As a proof-of-concept, we were able to port the MiniWoB benchmark with minimal effort (details in § B.2), with plans for additional benchmarks such as WebArena and WebShop.
@@ 5.1 Agent Design
We implement a simple web agent design with chain-of-thought prompting (Wei et al., 2022b), and we evaluate the performance of our agent across two axes: (1) the underlying LLM, and (2) the use of BrowserGym features.
@@ 5.2 Experimental Protocol
Standard Error:  To be able to run a range of experiments under a fixed budget, we limit the experiments to 10 seeds per task. After averaging results over the benchmark or subset of it, we usually observe a sufficiently low standard error to draw the needed conclusions. We use stratified bootstrap [^4] We sample with replacement for each task independently and we average across the tasks and seeds. to obtain 1,000 samples of the mean and report the mean and standard deviation of these means as success rate and standard error.
Max step:  For all tasks, we allow a maximum of 10 steps per episode. This ensures that a low-performing agent will not wander around for too long if it’s incapable of accomplishing the task; 10 steps are considered sufficient for all MiniWoB tasks. On the other hand, some WorkArena tasks can require a single-action agent more than 25 steps to accomplish, although 10 are enough in multi-action mode.
@@ 5.3 Baselines on WorkArena
Tab. 1 reports the zero-shot performance of these agents on two benchmarks: MiniWoB (Liu et al., 2018) and our proposed WorkArena. Performance is measured in terms of success rate. We emphasize our key findings below.
Table 1: Success rate±$\pm$±Standard error (SR $\pm$±SE) of all agents on MiniWoB and WorkArena. Bolded numbers represent the average success rate over the entire corresponding benchmark.
@@
CodeLLAMA’s Shortcomings in WorkArena:  Our analysis sheds light on why CodeLLAMA, despite its reasonable zero-shot performance on MiniWoB, fails to accomplish any task within WorkArena. We categorize and summarize the primary failure modes encountered, with details in § C. Initially, CodeLLAMA frequently replicates actions from prompt examples, which are unsuitable for a given context (§ C.1). Additionally, it misapplies both high-level (§ C.2) and Python APIs (§ C.3), and struggles with the intricacies of handling iFrames in Playwright (§ C.3). On the rare occasions that CodeLLAMA does execute relevant actions and navigates a WorkArena instance, it nevertheless fails to meet its objectives. These observations underscore a significant disparity between leading open-source LLMs and their closed counterparts further highlighting the rigorous challenge posed by our benchmark.
GPT-4’s Superiority in WorkArena:  The data unequivocally demonstrate GPT-4’s dominance over GPT-3.5 and CodeLlama within WorkArena. In the relatively simpler task category of menu navigation, GPT-4 achieves 95% success, in stark contrast to GPT-3.5’s 25% and CodeLlama’s 0%. The performance disparity between the LLMs is significantly more pronounced here than in MiniWoB, aligning with findings on the emergent properties of AI systems (Wei et al., 2022a). As the complexity of tasks increases, the necessity for a more advanced LLM to achieve any score becomes apparent, with noticeable improvements in performance correlating with the enhancement of capabilities.
@@ 5.5 Effect of BrowserGym Features
BrowserGym improves performance:  Tabs. 2 and 3 shows that adding observation- or action-space features often helps performance. While not always the case, we can observe a gap of more than 10% between the most basic configuration and the best one using GPT-3.5 on MiniWoB.
2D features do not help for WorkArena:  Various MiniWoB tasks require 2D understanding of the UI and interaction with $x,y$ coordinates; we observe a notable performance gain when introducing these 2D features in Tab. 2. On the other hand, none of them are useful on WorkArena since the proposed tasks can all be solved using bid only.
More is not always better:  As we gradually add more features, the prompt becomes longer and seems to overwhelm the LLM. This is mostly observed on WorkArena where most features suggest a performance drop. We hypothesise that since the accessibility tree is already large, adding more features is likely to distract the agent instead of helping it. We also note that in WorkArena, more features imply more accessibility tree truncation with GPT-3.5 and CodeLlama, due to limited context length (Fig. 4).
@@ A.1 Tasks
Lists (12 tasks) FilterAssetList
1,000
FilterChangeRequestList
1,000
FilterHardwareList
1,000
FilterIncidentList
1,000
FilterServiceCatalogItemList
1,000
FilterUserList
1,000
SortAssetList
1,000
SortChangeRequestList
1,000
SortHardwareList
1,000
SortIncidentList
1,000
SortServiceCatalogItemList
1,000
SortUserList
1,000
Forms (5 tasks) CreateChangeRequest
1,000
CreateIncident
1,000
CreateHardwareAsset
1,000
CreateProblem
1,000
CreateUser
1,000
Knowledge Bases (1 task) KnowledgeBaseSearch
1,000
Service Catalogs (9 tasks) OrderDeveloperLaptopMac
1,000
OrderIpadMini
80
OrderIpadPro
60
OrderSalesLaptop
1,000
OrderStandardLaptop
1,000
OrderAppleWatch
10
OrderAppleMacBookPro15
10
OrderDevelopmentLaptopPC
40
OrderLoanerLaptop
350
Menus (2 tasks) AllMenu
1,000
Impersonation
600
Total 23,150
@@ B.1 Action Space
bid fill(bid, text)
Fill an input field with text.
click(bid, button)
Click an element.
dblclick(bid, button)
Double-click an element.
hover(bid)
Hover the mouse over an element.
press(bid, key_comb)
Focus an element and press a combination of keys.
focus(bid)
Focus an element.
clear(bid)
Clear an input field.
select_option(bid, options)
Select one or multiple options in a drop-down element.
drag_and_drop(from_bid, to_bid)
Drag and drop one element to another.
coord mouse_move(x, y)
Move the mouse to a location.
mouse_down(x, y, button)
Move the mouse to a location then press and hold a mouse button.
mouse_up(x, y, button)
Move the mouse to a location then release a mouse button.
mouse_click(x, y, button)
Move the mouse to a location and click a mouse button.
mouse_dblclick(x, y, button)
Move the mouse to a location and double-click a mouse button.
mouse_drag_and_drop(from_x, from_y, to_x, to_y)
Drag and drop from a location to a location.
keyboard_down(key)
Press and holds a keyboard key.
keyboard_up(key)
Release a keyboard key.
keyboard_press(key_comb)
Press a combination of keys.
keyboard_type(text)
Types a string of text through the keyboard.
keyboard_insert_text(text)
Insert a string of text in the currently focused element.
tab new_tab()
Open a new tab.
tab_close()
Close the current tab.
tab_focus(index)
Bring a tab to front (activate tab).
nav go_back()
Navigate to the previous page in history.
go_forward()
Navigate to the next page in history.
goto(url)
Navigate to a url.
misc scroll(dx, dy)
Scroll pixels in X and/or Y direction.
send_msg_to_user(text)
Send a message to the user in the chat.
noop()
Do nothing.
python Any python code (unsafe!)
Executes code with playwright, the active page and the send_msg_to_user(text) primitive available.

### v2

@@ Abstract
Figure 1: Overview of contributions: (a) WorkArena is a benchmark of 29 web tasks and 18,050 unique instances that cover common ways of interacting with the ServiceNow Platform, a widely-used enterprise software platform. (b) BrowserGym is a Python environment for designing and evaluating web agents, which includes a rich set of actions and multimodal observations (shown here the HTML contents of the page, its accessibility tree, and the raw pixels after browser rendering).
@@ 1 Introduction
  • WorkArena: A realistic benchmark of enterprise-related tasks for web agents comprising 18,050 unique task instances (§ 3, Fig. 1a);
  • BrowserGym: A new framework for the development and evaluation of web agents, compatible with previous benchmarks like WebArena (Zhou et al., 2023), MiniWoB (Liu et al., 2018; Shi et al., 2017a) and WebShop (Yao et al., 2022), that offers a richer set of multimodal observations (e.g., screenshot, accessibility tree, screen coordinates), a broader set of actions (e.g., Python code and high-level primitives), and supports chat-based interactions (§ 4, Fig. 1b); Surprisingly, these features contribute to bringing our GPT-4 agent at the top of the leaderboard on WebArena, with a score of 25.4%, contrasting with the score of the original paper, 14.4%.
@@ 2 Related Works
Benchmarks for web agents:  Early benchmarks for web agents were based on synthetic web environments where agents were tasked with performing low-level keyboard and mouse actions (Shi et al., 2017b). Notable examples are MiniWoB (Shi et al., 2017a; Liu et al., 2018), which offer a collection of 125 toy web tasks ranging from clicking a specific button to using a basic text editor, and WebShop (Yao et al., 2022), a simulated e-commerce website with shopping tasks that require searching and browsing a catalog of items. More recently, Zhou et al. (2023) introduced WebArena, a collection of 190 tasks based on realistic websites that emulate real-world domains such as e-commerce, social forums, collaborative software development, and content management. WebArena is a notoriously challenging benchmark, with a success rate of 14% for a state-of-the-art web agent based on GPT-4, and 78% for human agents. Deng et al. (2023) proposed Min2Web, a large-scale dataset of 2,000 web interactions from 137 websites curated by human annotators. Similarly, Lù et al. (2024) propose WebLINX, a curated dataset of web interactions composed of 2337 expert demonstrations from 155 different real-world websites. In WebLINX, each task is composed of a turn-based chat dialogue averaging 43 interactions per task. Last, He et al. (2024) propose 300 information-retrieval tasks, from 15 real-world consumer websites (e.g., Amazon, Coursera, Booking), which are used to evaluate WebVoyager, a vision-based web agent.
@@ 3 WorkArena – An Enterprise Benchmark
WorkArena consists of a suite of 29 tasks and 18,050 unique instances that cover core interactions with the ServiceNow platform, such as navigating lists, filling forms, searching knowledge bases, utilizing service catalogs, and navigating via menus (see Fig. 1(a)). Collectively, these tasks are representative of a wide array of common operations that employees, like IT, administrative, and white-collar staff, perform on a daily basis.
@@ 3.1 WorkArena Tasks
Lists:  We consider 12 list-based tasks, which can be grouped into two categories: filtering and sorting. The former consists of using the UI to construct a complex filter with 1 to 5 conditions. The latter consists of using the UI to sort the list based on up to 3 columns. In both cases, the interaction with the UI is non-trivial and requires opening a hidden menu, adding the right number of conditions, and filling them out accordingly. There are 6 tasks of each type, each corresponding to a different data table (e.g., users, incidents). In both cases, client-side validation is used to verify that the resulting list satisfies the expected conditions. Together, these tasks yield 6,900 instances.
@@ 4.2 An Ideal Experimental Framework
Extensibility:  BrowserGym is easily extensible to additional benchmarks. We currently support MiniWoB § B.2), WebArena (Zhou et al., 2023) and WorkArena. We encourage the community to add new benchmarks or new agents to this platform. We also offer support to reduce the friction of adding new benchmarks to the platform. Please reach out on GitHub for assistance.
@@ 5.1 Agent Design
We implement a simple web agent with chain-of-thought prompting (Wei et al., 2022b), and we evaluate its performance across two axes: (1) the underlying LLM, and (2) the use of BrowserGym features.
@@ 5.2 Experimental Protocol
Standard Error:  To be able to run a range of experiments under a fixed budget, we limit the number of seeds to 10 per task. After averaging results over each benchmark (or subset of it), we usually observe a sufficiently low standard error to draw the needed conclusions. We use stratified bootstrap [^4] We sample with replacement for each task independently and we average across the tasks and seeds. to obtain 1,000 samples of the mean and report the average and standard deviation of these means as success rate and standard error.
Max step:  For all tasks, we allow a maximum of 10 steps per episode. This ensures that a low-performing agent will not wander around for too long if it’s incapable of solving the task; 10 steps are considered sufficient for all MiniWoB tasks. On the other hand, some WorkArena tasks can require a single-action agent more than 25 steps to accomplish, although 10 are enough in multi-action mode.
@@ 5.3 Baselines on WorkArena
Tab. 1 reports the zero-shot performance of these agents on three benchmarks: MiniWoB (Liu et al., 2018), WebArena (Zhou et al., 2023) and our proposed WorkArena. Performance is measured in terms of success rate. We emphasize our key findings below.
Table 1: Success rate±$\pm$±Standard error (SR $\pm$±SE) of all agents on MiniWoB, WorkArena and WebArena. Bolded numbers represent the average success rate over the entire corresponding benchmark.
[…]
@@
WebArena (812 tasks) 25.4 $\pm$0.6 13.1 $\pm$0.6 0 $\pm$0
   Content-and-config (411 tasks) 28.7 $\pm$0.9 21.9 $\pm$0.9 0 $\pm$0
   Information-seeking (325 tasks) 23.1 $\pm$1.0 4.0 $\pm$1.0 0 $\pm$0
   Navigation (76 tasks) 17.1 $\pm$2.1 3.9 $\pm$2.0 0 $\pm$0
@@
CodeLLAMA’s Shortcomings:  Our analysis sheds light on why CodeLLAMA, despite its reasonable zero-shot performance on MiniWoB, fails to accomplish any task within WorkArena. We categorize and summarize the primary failure modes encountered, with details in § C. Initially, CodeLLAMA frequently replicates actions from prompt examples, which are unsuitable for a given context (§ C.1). Additionally, it misapplies both high-level (§ C.2) and Python APIs (§ C.3), and struggles with the intricacies of handling iFrames in Playwright (§ C.3). On the rare occasions that CodeLLAMA does execute relevant actions and navigates a WorkArena instance, it nevertheless fails to meet its objectives. These observations underscore a significant disparity between leading open-source LLMs and their closed counterparts further highlighting the rigorous challenge posed by our benchmark.
GPT-4’s Superiority:  The data unequivocally demonstrate GPT-4’s dominance over GPT-3.5 and CodeLlama within WorkArena. In the relatively simpler task category of menu navigation, GPT-4 achieves 95% success, in stark contrast to GPT-3.5’s 25% and CodeLlama’s 0%. The performance disparity between the LLMs is significantly more pronounced here than in MiniWoB and WebArena, aligning with findings on the emergent properties of AI systems (Wei et al., 2022a). As the complexity of tasks increases, the necessity for a more advanced LLM to achieve any score becomes apparent, with noticeable improvements in performance correlating with the enhancement of capabilities.
@@ 5.5 Effect of BrowserGym Features
BrowserGym improves performance:  Tabs. 2 and 3 shows that adding observation- or action-space features often helps performance. While not always the case, we can observe a gap of more than 10% between the most basic configuration and the best one using GPT-3.5 on MiniWoB. Moreover, our GPT4 agents obtains state-of-the-art performance on WebArena with 25.4% success rate.
2D features do not help for WorkArena:  Several MiniWoB tasks require 2D understanding of the UI and interaction with $x,y$ coordinates; we observe a notable performance gain when introducing 2D features in Tab. 2. On the other hand, on WorkArena the tasks seem best solved using bid only.
More is not always better:  As we gradually add more features, the prompt becomes longer and seems to overwhelm the LLM. This is mostly observed on WorkArena where additional features lead to a performance drop. We hypothesise that since the AXTree is already large, adding more features is likely to distract the agent instead of helping it. We also note that in WorkArena, more features imply more prompt truncation with GPT-3.5 and CodeLlama, due to limited context length (Fig. 4).
@@ A.1 Tasks
Lists (12 tasks) FilterAssetList 1,000
FilterChangeRequestList 1,000
FilterHardwareList 1,000
FilterIncidentList 1,000
FilterServiceCatalogItemList 1,000
FilterUserList 1,000
SortAssetList 150
SortChangeRequestList 150
SortHardwareList 150
SortIncidentList 150
SortServiceCatalogItemList 150
SortUserList 150
Forms (5 tasks) CreateChangeRequest 1,000
CreateIncident 1,000
CreateHardwareAsset 1,000
CreateProblem 1,000
CreateUser 1,000
Knowledge Bases (1 task) KnowledgeBaseSearch 1,000
Service Catalogs (9 tasks) OrderDeveloperLaptopMac 1,000
OrderIpadMini 80
OrderIpadPro 60
OrderSalesLaptop 1,000
OrderStandardLaptop 1,000
OrderAppleWatch 10
OrderAppleMacBookPro15 10
OrderDevelopmentLaptopPC 40
OrderLoanerLaptop 350
Menus (2 tasks) AllMenu 1,000
Impersonation 600
Total 18,050
@@ B.1 Action Space
bid fill(bid, text) Fill an input field with text.
click(bid, button) Click an element.
dblclick(bid, button) Double-click an element.
hover(bid) Hover the mouse over an element.
press(bid, key_comb) Focus an element and press a combination of keys.
focus(bid) Focus an element.
clear(bid) Clear an input field.
select_option(bid, options) Select one or multiple options in a drop-down element.
drag_and_drop(from_bid, to_bid) Drag and drop one element to another.
coord mouse_move(x, y) Move the mouse to a location.
mouse_down(x, y, button) Move the mouse to a location then press and hold a mouse button.
mouse_up(x, y, button) Move the mouse to a location then release a mouse button.
mouse_click(x, y, button) Move the mouse to a location and click a mouse button.
mouse_dblclick(x, y, button) Move the mouse to a location and double-click a mouse button.
mouse_drag_and_drop(from_x, from_y, to_x, to_y) Drag and drop from a location to a location.
keyboard_down(key) Press and holds a keyboard key.
keyboard_up(key) Release a keyboard key.
keyboard_press(key_comb) Press a combination of keys.
keyboard_type(text) Types a string of text through the keyboard.
keyboard_insert_text(text) Insert a string of text in the currently focused element.
tab new_tab() Open a new tab.
tab_close() Close the current tab.
tab_focus(index) Bring a tab to front (activate tab).
nav go_back() Navigate to the previous page in history.
go_forward() Navigate to the next page in history.
goto(url) Navigate to a url.
misc scroll(dx, dy) Scroll pixels in X and/or Y direction.
send_msg_to_user(text) Send a message to the user in the chat.
noop() Do nothing.
python Any python code (unsafe!) Executes code with playwright, the active page and the send_msg_to_user(text) primitive available.